A gas has a pressure of 630 mmHg and at 300 K. At what temperature would the pressure be 101.3 mmHg if the volume remains constant?
1
Expert's answer
2020-10-14T05:53:57-0400

for ideal gas At constant volume, pressure is proportional to the temperature ,P1/T1=P2/T2

Or,T2=P2T1/P1,where P represent the pressure,and T represent the temperature,1 for initial and 2 for final,

T2=(101.3*300)/630=48.23K

Hence at temperature 48.23K the pressure will be 101.3mmHg